Simplify the following Boolean functions using four-variable maps:

  1. F(w, x, y, z) = Σ (1, 4, 5, 6, 12, 14, 15)
  2. F(w, x, y, z) = Π (0, 1, 4, 5, 6, 7, 8, 9)
  3. AB’C + B’C’D’ + BCD + ACD’ + A’B’C+ A’BC’D
  4. (A xor B)’ (C xor D)
Answer #1

Solution: a flw,x,y,z) = {(1,4,5,6,12, 14, 15) K-Map? W Y Z yz YZ YZ YZ four bairs are formed wx V wxIA wxil oL 1]0% F(d). f (A,B,CD) = (AM) B) (COD) = (AB + AB) Ced+(0) = (ABJ (A B) (C+C)(C+D)(C+) (0+D) = (A+6) (4+B) (C+D) (070